# To explain the muon excess phenomenon in cosmic rays based on the gluon condensation model

### v1

@@ Abstract
Ultrahigh-energy cosmic rays are often characterized indirectly by analyzing the properties of secondary cosmic ray particles produced in the collisions with air nuclei, the particle number $N_{\mu}$ of muon and the depth of shower maximum $X_{\mathrm{max}}$ after air shower cascade are mostly studied to infer the energy and mass of the incident cosmic rays. While, researches have shown that there is a significant excess comparing the observed number of muons arriving at the ground from extensive air showers (EAS) with the simulations by using the existing cosmic ray hadronic interaction model. To explain this muon excess phenomenon, a new theoretical model, the gluon condensation model (GC model), is introduced in this paper and simulated by using the AIRES engine. We asumme that the GC effect appearing mainly in the first colliding of the cascade, leads to a significant increase in the strangeness production, accodingly, the production rate of kaons is improved appearantly. The model assumes that only pions and kaons are the new prodcutions in the hadron cascades. It is found that, considering the GC effect, the value of $n_{K}/n_{\pi}$ increases and more energy of the incident cosmic rays paticipate in hadron cascades, and then increase the number of mouns finally. This model provides a new theoretical possibility to explain the muon excess puzzle.
@@
The number of muons $N_{\mu}$ and depth of shower maximum $X_{\mathrm{max}}$ are the key parameters for studying the cosmic rays, which are sensitive to several properties of hadronic interactions. The mean $N_{\mu}$ is mainly sensitive to the multiplicity, the energy fraction of $\pi^{0}$ (which carries the proportion of the incident energy in hadronic interactions and then turns to electromagnetic interactions), and the primary mass of cosmic rays. The mean $X_{\mathrm{max}}$ is mainly sensitive to the cross-section, elasticity, multiplicity, and primary mass Allen & Farrar (2013), and the measurement of $X_{\mathrm{max}}$ shows well agreement with the model estimates. It can be seen that only by altering the energy fraction of $\pi^{0}$ which can increase $N_{\mu}$ while without affecting the $X_{\mathrm{max}}$. To address the muon excess issue, the ”fireball model” proposed by Anchordoqui et al. (2017) enhances the strangeness to suppress $\pi^{0}$ production. Additionally, in Anchordoqui et al. (2022), a model is proposed where swapping $\pi$ and $K$ particles is used to explain the muon excess, with $f_{s}$ representing the swapping probability of $\pi$ to $K$. The work of Baur et al. (2023) employs a simplified core-corona method to estimate the muon production when statistical hadronization is present. The core part corresponds to the normal particles produced by hadronic interactions, while the corona part corresponds to the production of ”statistical” or ”thermal” particles, each with its own weight summing up to 1.
@@
The aim of this paper is to elucidate the muon excess problem in cosmic rays using the gluon condensation (GC) model. Air shower cascades are driven by collisions between ultrahigh energy primary cosmic rays with the hadrons in atmosphere. During the collisions, the parton distribution in small $x$ region is dominated by gluons. The evolution of gluons follows the equations of quantum chromodynamics (QCD), but in the small $x$ region, traditional evolution equations should be modified due to the recombination among partons. The modified equation, known as the ZSR equation Zhu & Ruan (1999); Zhu & Shen (2004), predict that gluons in hadrons may converge at a critical momentum $k_{c}$ in very high energy region, which is called the gluon condensation (GC). By simulation and analysis of the gluon condensation effects in collision processes, we found that the GC effect significantly increases the production of pions and kaons, especially the production rate of $n_{K}/n_{\pi}$ being much higher than that in normal collidings Wong (1994). Simulation results based on the GC model show a substantial increase in the number of muons, the problem of muon excess may be solved.
@@ 2.1 QCD dynamic evolution equation
In the infinite frame of momentum, hadrons are composed of partons, namely quarks and gluons. Partons evolve according to the evolution equations based on Quantum Chromodynamics (QCD) within the framework of the standard model. DGLAP (Dokshitzer-Gribov-Lipatov-Altarelli-Parisi) equationGribov & Lipatov (1972); Dokshitzer (1977); Altarelli & Parisi (1977) and BFKL (Balitsky-Fadin-Kuraev-Lipatov) equationFadin et al. (1975); Kuraev et al. (1976); Balitsky & Lipatov (1978) are the standard parton evolution equations. In small $x$ region, BFKL equation is more suitable since the parton evolution in it includes the information of transverse momentum. The BFKL equation reads
$-x\frac{\partial F(x,k_{\perp}^{2})}{\partial x}=\frac{3\alpha_{s}k_{\perp}^{2}}{\pi}\int_{k_{0}^{2}}^{\infty}\frac{dk_{\perp}^{\prime 2}}{k_{\perp}^{\prime 2}}$$\left\{\frac{F(x,k_{\perp}^{\prime 2})-F(x,k_{\perp}^{2})}{|k_{\perp}^{\prime 2}-k_{\perp}^{2}|}\right.$ (1)
$\left.+\frac{F(x,k_{\perp}^{2})}{\sqrt{k_{\perp}^{4}+4k_{\perp}^{\prime 4}}}\right\},$
where the function $F(x,k_{\perp}^{2})$ is the unintegrated gluon distribution function. In works Zhu & Ruan (1999); Zhu & Shen (2004), W.Zhu etc., took into account the recombination effects among gluons, using the Time-Ordered Perturbative Theory (TOPT), they obtained a nonlinear dynamical evolution equation, named as the ZSR equation. The equation is
@@
$-\frac{81}{16}\frac{\alpha_{s}^{2}}{\pi R_{N}^{2}}\int_{k_{0}^{2}}^{\infty}\frac{dk_{\perp}^{\prime{2}}}{k_{\perp}^{\prime{2}}}\left\{\frac{k_{\perp}^{2}F^{2}(x,k_{\perp}^{\prime{2}})-k_{\perp}^{\prime{2}}F^{2}(x,k_{\perp}^{2})}{k_{\perp}^{\prime{2}}|k_{\perp}^{\prime{2}}-k_{\perp}^{2}|}\right.$
$\qquad\qquad\qquad\qquad\qquad\left.+\frac{F^{2}(x,k_{\perp}^{2})}{\sqrt{k_{\perp}^{4}+4k_{\perp}^{\prime{4}}}}\right\}.$
The last term in the equation is the additional terms which includes the gluon recombination effects. The specific nonlinear terms may couple stochastic perturbations, leading to chaotic phenomena. One criterion for the appearance of chaotic phenomena is the presence of a positive Lyapunov exponent Wei et al. (2008); Zhu et al. (2016); Zhu & Lan (2017), indicating the system high sensitivity to small changes in initial conditions. The chaos alone cannot produce the GC solutions in the ZSR equation. The TOPT regularized nonlinear kernel plays its second important role for converting the chaotic vibration, which occurs in the evolution equation when the Bjorken variable $x$ is less than the critical point $x_{c}$, into the strong shadowing and antishadowing effect. The GC effect results from the combined influence of chaos and antishadowing effect. Potential perturbation arises chaotic oscillations in the BFKL dynamics, generating the positive antishadowing corrections, which then convert the shadowed gluons to the gluons with the critical momentum $\left(x_{c},k_{c}\right)$ Zhu et al. (2022). The gluon condensation is a natural consequence of gluons converging to the critical momentum state. The specific distribution function of gluons can be approximately represented as a delta function, as illustrated in Figure 1.
@@ 2.2 Particle generation in GC state
To deeply understand the hadronization process in QCD, a recent research by Roberts et al. based on the Dyson-Schwinger (DS) equations showed that the effective quark mass in hadrons might be dynamically generated due to the gluon nonlinearity Mathieu & Vento (2010); Chang et al. (2014); Roberts (2020); Rodríguez-Quintero et al. (2020). In this mechanism the nonperturbative gluon propagator is constructed by lots of partons (i.e., gluons and sea quarks) in the hadron infinite momentum frame. Massless quark gains a small current mass through the Higgs mechanism and then further acquires the effective mass through the nonperturbative gluon propagator. The gluon condensates, which are rich in gluons, greatly generate the number of the constituent quarks and increase the number of pions and kaons. As a limit, it is imaged that at the center-of-mass system almost all available collision energy is used to produce pions and kaons. In this paper, the gluon condensation model is exploited to calculate the number of secondary particles of proton-proton collisions in the high-energy cosmic ray. The AIRES simulation engine Sciutto (1999), which provides full space-time particle propagation in a realistic environment, is adopted to describe the shower evolution.
@@
where $E_{pi}^{*}$ is the energy of the leading proton at the center-of-mass system, $\gamma_{i}$ is the corresponding Lorentz factor, $n_{\pi}$ and $n_{K}$ are the number of pions and kaons produced after the collision, respectively. In this work, we assume that the secondary particles resulting from the collisions between the incident protons with the proton targets, specifically including leading protons, pions, and kaons. These secondary particles ( protons, pions, and kaons) continue to collide with protons in the atmosphere until they reach a critical energy and decay into muons. Using the inelasticity $K\sim 0.5$ Gaisser et al. (2016)and setting
@@ 3.1 Mechanism of muon production
Extensive air showers (EAS) develop in complex processes as combination of electromagnetic cascades and hadronic cascades. It is necessary to perform detailed numerical simulations of air showers to infer the properties of the primary cosmic rays that initiate them. Thus, Heitler presented a simple model of electromagnetic(EM) cascade development Heitler (1984). The electromagnetic cascade involves two physical processes: the $e^{\pm}$ pairs production by photons and the bremsstrahlung by $e^{\pm}$. As depicted in the top panel of Figure 2, high energy photons interact with atmospheric particles to produce $e^{\pm}$ pairs. The $e^{\pm}$ pairs continue to interact with nuclei during the propagation to produce bremsstrahlung, which generate photons. This cascade of photons and electrons along the shower axis is referred to as the electromagnetic cascade. Matthews generalised this concept of the Heitler model to hadronic interactions Matthews (2005). As sketched in the bottom panel of Figure 2, the hadronic cascade in basic Heitler-Matthews model is approximated by a pure pion shower. Cosmic ray protons with an primary energy of $E_{0}$ collide with nuclei in the atmosphere, producing secondary high-energy hadrons. Neutral pions, with short lifetimes, decay into photons, leading to electromagnetic cascades. Charged pions continue to interact with particles in the atmosphere, producing hadronic interactions. Through a fixed amount of slant lengths, more $\pi^{\pm}$ and $\pi^{0}$ are produced until the energy of $\pi^{\pm}$ reach the critical energy $E_{c}^{\pi^{\pm}}\sim 115\mathrm{GeV}$ Thunman et al. (1996). At this point, $\pi^{\pm}$ no longer undergo interactions, and the hadronic cascade stops. Then, all charged pions decay into muons at a 1:1 ratio. Expect pions, kaons can also decay into muons, with the critical energies $E_{c}^{K^{\pm}}\sim 850\mathrm{GeV}$, $E_{c}^{K^{0}_{L}}\sim 210\mathrm{GeV}$, and $E_{c}^{K^{0}_{S}}\sim 30\mathrm{TeV}$ Thunman et al. (1996). The decay processes of $\pi$ and $K$ are shown in Eq.8.
Figure 2: Electromagnetic cascade(top panel). Hadronic cascadeUlrich et al. (2011)(bottom panel)
@@
Thus, the number of muons corresponds to the number of charged pions and kaons at the end of the hadronic cascade. In the basic Heitler-Matthews model, setting up in the first interaction produces a fixed number $n_{\mathrm{mult}}$ of pions, among which the fraction of charged pions is $r=2/3$ Albrecht et al. (2022). Thus, after undergoing $n_{c}$ critical cascade interactions, the total number of charged pions $N_{\mathrm{ch}}$, which is equal to the number of muons $N_{\mu}$, is given by $N_{\mu}=N_{\mathrm{ch}}=(rn_{\mathrm{mult}})^{n_{c}}$. When the energy of the charged pions equals the critical energy $E_{c}$, the cascade reaches the critical cascade generation $n_{c}$, $N_{\mu}$ can be caclulated as:
@@
where $E_{0}$ is the primary energy of cosmic rays, and $\beta=\frac{\ln(rn_{\mathrm{mult}})}{\ln n_{\mathrm{mult}}}$. According to the Heitler-Matthews model, each generation of particles includes $\pi^{0}$ carrying one-third of the energy into the electromagnetic cascade. Taking into account the hadronic interactions and decay of kaons, as well as some other hadronic resonances, approximately $25\%$ of the energy in each generation of particles is transferred to the electromagnetic cascade Thunman et al. (1996). Finally, the electromagnetic cascade consumes about $90\%$ of the primary particle energy, while the remaining $10\%$ is carried by muons and neutrinos. This suggests a method to increase the number of muons by reducing the proportion of energy in the electromagnetic cascade and increasing the proportion in the hadronic cascade. Therefore, it is necessary to focus on reducing the production of $\pi^{0}$. Kaons (such as $K^{+}$, $K^{-}$, $K^{0}_{S}$, $K^{0}_{L}$) participate in hadronc interactions, as in Eq.8, after reaching the critical energy, their decay products include not only muons but also pions. The branching ratio of kaon decaying into $\pi^{0}$ is much lower than decaying into muon, as reported in Patrignani et al. (2016). Thus, in this view, increasing the production of secondary particle kaons can increase the number of muons in the simulation after p-p collisions. Since kaons consist of a strange quark and a light quark, increasing the production of strange quarks is a key factor considered in this study.

### v2

# Explaining muon excess in cosmic rays using the gluon condensation model

@@ Abstract
Ultrahigh-energy cosmic rays are often characterized indirectly by analyzing the properties of secondary cosmic ray particles produced in the collisions with air nuclei. The particle number $N_{\mu}$ of muon and the depth of shower maximum $X_{\mathrm{max}}$ after air shower cascade are mostly studied to infer the energy and mass of the incident cosmic rays. Research have shown that there is a significant excess in the observed number of muons arriving at the ground from extensive air showers (EAS) compared to the simulations using the existing cosmic ray hadronic interaction model. To explain this muon excess phenomenon, a new theoretical model, the gluon condensation model (GC model), is introduced in this paper and simulated by using the AIRES engine. We assume that the GC effect appears mainly in the first collision of the cascade leading to a significant increase in the strangeness production, consequently, the production rate of kaons is increased and $n_{K}/n_{\pi}$ is greater than the value of the usual hadronic interaction process. In the calculation, the model assumes that only pions and kaons are produced in GC state. The increase of strange particle yield would mean that the energy transferred from the hadronic cascade to electromagnetic cascade through $\pi^{0}\rightarrow 2\gamma$ decay is reduced. This would in turn increase the number of muons at the ground level due to meson decays. Our model provides a new theoretical possibility to explain the muon excess puzzle.
Cosmic ray showers(327) — Particle astrophysics(96) — High energy astrophysics(739)
@@ 1 Introduction
The number of muons $N_{\mu}$ and depth of shower maximum $X_{\mathrm{max}}$ are the key parameters for studying the cosmic rays, which are sensitive to several properties of hadronic interactions. The mean $N_{\mu}$ is mainly sensitive to the multiplicity, the energy fraction of $\pi^{0}$ (which carries the proportion of the incident energy in hadronic interactions and then transfers to electromagnetic interactions), and the primary mass of cosmic rays. The mean $X_{\mathrm{max}}$ is mainly sensitive to the cross-section, elasticity, multiplicity, and primary mass Allen & Farrar (2013), and the measurement of $X_{\mathrm{max}}$ shows well agreement with the model estimates. It can be seen that only by altering the energy fraction of $\pi^{0}$, $N_{\mu}$ can be increased while $X_{\mathrm{max}}$ is kept unchanged. To address the muon excess issue, the ”fireball model” proposed by Anchordoqui et al. (2017) enhances the strangeness to suppress $\pi^{0}$ production. Additionally, in Anchordoqui et al. (2022), a model is proposed where swapping $\pi$ and $K$ particles is used to explain the muon excess, with $f_{s}$ representing the swapping probability of $\pi$ to $K$. The work of Baur et al. (2023) employs a simplified core-corona method to estimate the muon production when statistical hadronization is present. The core part corresponds to the normal particles produced by hadronic interactions, while the corona part corresponds to the production of ”statistical” or ”thermal” particles, each with its own weight summing up to 1. Moreover, a novel studyAab et al. (2021) on the shower-to-shower fluctuations of $N_{\mu}$ by Pierre Auger Observatory suggests that this muon deficit in models might be a small deficit at every stage of the shower that accumulates along the shower development, and they also mentioned that a very particular modification of the first interaction may also change $N_{\mu}$ without changing the fluctuations. Although the research about muon puzzle are becoming increasingly detailed, the truth remains elusive, necessitating the further investigations.
Regarding the muon excess issue, we tend to believe that it’s the initial collision in air showers which cause the main difference between the measurements and model predictions. In ultrahigh-energy air showers, the interaction energy of their first collision may significantly exceed the energy of accelerators, potentially leading to new physical phenomena. Zhu et al., Zhu & Ruan (1999); Zhu & Shen (2004) have proposed that in ultrahigh-energy interactions, the gluon distribution in hadrons may undergo condensation, which would influence the hadronization products and subsequently the final muon production. The aim of this paper is to use the gluon condensation model in the first collision of the cascades to resolve the muon excess problem.
@@
Air shower cascades are driven by collisions between ultrahigh energy primary cosmic rays with the hadrons in atmosphere. During the collisions, the parton distribution in small $x$ region is dominated by gluons, where $x$ is the parton momentum fraction (Bjorken-$x$). The evolution of gluons follows the equations of quantum chromodynamics (QCD), but in the small $x$ region, traditional evolution equations should be modified due to the recombination among partons. The modified equation, known as the Zhu-Shen-Ruan (ZSR) equation Zhu & Ruan (1999); Zhu & Shen (2004), predict that gluons in hadrons may converge at a critical momentum $k_{c}$ in very high energy region, which is called the gluon condensation (GC). By simulation and analysis of the gluon condensation effects in collision processes, we found that the GC effect significantly increases the production of pions and kaons, especially the production rate of $n_{K}/n_{\pi}$ being much higher than that in normal collidings Wong (1994). The increase of kaon yield would mean that the energy transferred from the hadronic cascade to electromagnetic cascade through $\pi^{0}\rightarrow 2\gamma$ decay is reduced. This would in turn increase the number of muons at the ground level due to meson decays, which likely resolve the problem of muon excess.
@@ 2.1 QCD dynamic evolution equation
In the infinite frame of momentum, hadrons are composed of partons, namely quarks and gluons. Partons evolve according to the evolution equations based on Quantum Chromodynamics (QCD) within the framework of the standard model. DGLAP (Dokshitzer-Gribov-Lipatov-Altarelli-Parisi) equationGribov & Lipatov (1972); Dokshitzer (1977); Altarelli & Parisi (1977) and BFKL (Balitsky-Fadin-Kuraev-Lipatov) equationFadin et al. (1975); Kuraev et al. (1976); Balitsky & Lipatov (1978) are the standard parton evolution equations. In small $x$ region, BFKL equation is more suitable since the parton evolution in it includes the information of transverse momentum, and the transverse momentum is one of the key parameters when considering the parton recombinations. The BFKL equation readsEnberg & Peschanski (2006)
$-x\frac{\partial F(x,k_{\perp}^{2})}{\partial x}=\frac{3\alpha_{s}k_{\perp}^{2}}{\pi}\int_{k_{0}^{2}}^{\infty}\frac{dk_{\perp}^{\prime 2}}{k_{\perp}^{\prime 2}}\left\{\frac{F(x,k_{\perp}^{\prime 2})-F(x,k_{\perp}^{2})}{|k_{\perp}^{\prime 2}-k_{\perp}^{2}|}+\frac{F(x,k_{\perp}^{2})}{\sqrt{k_{\perp}^{4}+4k_{\perp}^{\prime 4}}}\right\},$ (1)
where the function $F(x,k_{\perp}^{2})$ is the unintegrated gluon distribution function and $k_{\perp}$ is its transverse momentum, $\alpha_{s}$ is the effective QCD coupling constant. The BFKL equation describes the parton evolution process of one parton (with larger $x$) splitting into two partons (with smaller $x$). So, as $x$ decreases, the parton density increases, and gradually, the interactions between partons become non-negligible. The ZSR equation is a modified-BFKL equation which considered the parton recombination effects of twist-4 processes using the Time-Ordered Perturbative Theory(TOPT) Zhu & Ruan (1999); Zhu & Shen (2004). The ZSR equation is
@@
$-\frac{81}{16}\frac{\alpha_{s}^{2}}{\pi R_{N}^{2}}\int_{k_{0}^{2}}^{\infty}\frac{dk_{\perp}^{\prime{2}}}{k_{\perp}^{\prime{2}}}\left\{\frac{k_{\perp}^{2}F^{2}(x,k_{\perp}^{\prime{2}})-k_{\perp}^{\prime{2}}F^{2}(x,k_{\perp}^{2})}{k_{\perp}^{\prime{2}}|k_{\perp}^{\prime{2}}-k_{\perp}^{2}|}+\frac{F^{2}(x,k_{\perp}^{2})}{\sqrt{k_{\perp}^{4}+4k_{\perp}^{\prime{4}}}}\right\}.$
The last term in the above equation is the contribution of gluon recombination twist-4 Feynman diagrams, where $R_{N}$ is the effective radius of nucleon. This specific nonlinear term may couple stochastic perturbations, introducing complexity and instability into the system, magnifying small changes in initial conditions, and leading to chaotic phenomena. One criterion for the appearance of chaotic phenomena is the presence of a positive Lyapunov exponent Zhu et al. (2008, 2016); Zhu & Lan (2017), indicating the system high sensitivity to small changes in initial conditions. However, the chaos alone cannot produce the GC solutions in the ZSR equation. The TOPT regularized nonlinear kernel plays its second important role for converting the chaotic vibration, which occurs in the evolution equation when the Bjorken variable $x$ is less than the critical point $x_{c}$, into the strong shadowing and antishadowing effect. The GC effect results from the combined influence of chaos and antishadowing effect. Potential perturbation lead to chaotic oscillations in the BFKL dynamics, generating the positive antishadowing corrections, which then convert the shadowed gluons to the gluons with the critical momentum $\left(x_{c},k_{c}\right)$ Zhu et al. (2022). The gluon condensation is a natural consequence of gluons converging to the critical momentum state. The specific distribution function of gluons can be approximately represented as a delta function, as illustrated in Figure 1.
@@ 2.2 Particle generation in GC state
To deeply understand the hadronization process in QCD, a recent research by Roberts Roberts (2020) based on the Dyson-Schwinger (DS) equations showed that the effective quark mass in hadrons might be dynamically generated due to the gluon nonlinearity Mathieu & Vento (2010); Chang et al. (2014); Rodríguez-Quintero et al. (2020). In this mechanism the nonperturbative gluon propagator is constructed by lots of partons (i.e., gluons and sea quarks) in the hadron infinite momentum frame. Massless quark gains a small current mass through the Higgs mechanism and then further acquires the effective mass through the nonperturbative gluon propagator. In high-energy proton-proton collisions, the produced asymptotically free quarks and gluons cannot exist independently due to color confinement. These quarks and gluons interact by exchanging gluons, forming stable bound states known as hadrons. The gluon condensates, which are rich in gluons, greatly generate the number of the constituent quarks, thereby making the hadronization process more effective and increasing the number of pions and kaons. As a limit, it is imaged that at the center-of-mass system almost all available collision energy is used to produce pions and kaons. In this paper, the gluon condensation model is exploited to calculate the number of secondary particles of proton-proton collisions in the high-energy cosmic ray. The AIRES simulation engine Sciutto (1999), which provides full space-time particle propagation in a realistic environment, is adopted to describe the shower evolution.
@@
where $E_{pi}^{*}$ is the energy of the leading proton at the center-of-mass system, $\gamma_{i}$ is the corresponding Lorentz factor, $n_{\pi}$ and $n_{K}$ are the number of pions and kaons produced after the collision, respectively. In this work, we assume that the secondary particles resulting from the collisions between the incident protons with the proton targets, specifically including leading protons, pions, and kaons. These secondary particles (protons, pions, and kaons) continue to collide with protons in the atmosphere until they reach a critical energy, where particle decay becomes the main process. Using the inelasticity $K\sim 0.5$ Gaisser et al. (2016), we get
@@ 3.1 Mechanism of muon production
Extensive air showers (EAS) develop in complex processes as combination of electromagnetic cascades and hadronic cascades. It is necessary to perform detailed numerical simulations of air showers to infer the properties of the primary cosmic rays that initiate them. Thus, Heitler presented a simple model of electromagnetic(EM) cascade development Heitler (1984). The electromagnetic cascade involves two physical processes: the $e^{\pm}$ pairs production by photons and the bremsstrahlung by $e^{\pm}$. As depicted in the left panel of Figure 2, high energy photons interact with atmospheric particles to produce $e^{\pm}$ pairs. The $e^{\pm}$ pairs continue to interact with nuclei during the propagation to produce bremsstrahlung, which generate photons. This cascade of photons and electrons along the shower axis is referred to as the electromagnetic cascade. Matthews generalised this concept of the Heitler model to hadronic interactions Matthews (2005). As sketched in the right panel of Figure 2, the hadronic cascade in basic Heitler-Matthews model is approximated by a pure pion shower. Cosmic ray protons with an primary energy of $E_{0}$ collide with nuclei in the atmosphere, producing secondary high-energy hadrons. Neutral pions, with short lifetimes, decay into photons, leading to electromagnetic cascades. Charged pions continue to interact with particles in the atmosphere, producing hadronic interactions. Through a fixed amount of slant lengths, more $\pi^{\pm}$ and $\pi^{0}$ are produced until the energy of $\pi^{\pm}$ reach the critical energy $E_{c}^{\pi^{\pm}}\sim 115\mathrm{GeV}$ Thunman et al. (1996). At this point, $\pi^{\pm}$ no longer undergo interactions, and the hadronic cascade stops. Then, all charged pions decay into muons at a 1:1 ratio. Expect pions, kaons can also decay into muons, with the critical energies $E_{c}^{K^{\pm}}\sim 850\mathrm{GeV}$, $E_{c}^{K^{0}_{L}}\sim 210\mathrm{GeV}$, and $E_{c}^{K^{0}_{S}}\sim 30\mathrm{TeV}$ Thunman et al. (1996). The decay processes of $\pi$ and $K$ are shown in Eq.8.
Figure 2: Simplified casecade model of air shower. The left panel shows the electromagnetic cascade and the right panel is the hadronic cascade from Ulrich et al. (2011). In the hadronic cascade, solid (dashed) lines represent charged (neutral) pions.
@@
Thus, the number of muons corresponds to the number of charged pions and kaons at the end of the hadronic cascade. In the basic Heitler-Matthews model, assuming the production number of pions up in each interaction is $n_{\mathrm{mult}}$, among which the fraction of charged pions is $r=2/3$ Albrecht et al. (2022). Thus, after undergoing $n_{c}$ critical cascade interactions, the total number of charged pions $N_{\mathrm{ch}}$, which is equal to the number of muons $N_{\mu}$, is given by $N_{\mu}=N_{\mathrm{ch}}=(rn_{\mathrm{mult}})^{n_{c}}$. When the energy of the charged pions equals the critical energy $E_{c}$, the cascade reaches the critical cascade generation $n_{c}$, $N_{\mu}$ can be caclulated as:
@@
where $E_{0}$ is the primary energy of cosmic rays, and $\beta=\frac{\ln(rn_{\mathrm{mult}})}{\ln n_{\mathrm{mult}}}$. According to the Heitler-Matthews model, each generation of particles includes $\pi^{0}$ carrying one-third of the energy into the electromagnetic cascade. Taking into account the hadronic interactions and decay of kaons, as well as some other hadronic resonances, approximately $25\%$ of the energy in each generation of particles is transferred to the electromagnetic cascade Thunman et al. (1996). Finally, the electromagnetic cascade consumes about $90\%$ of the primary particle energy, while the remaining $10\%$ is carried by muons and neutrinos.
The production of muon is a result of the secondary particles decay from hadronic cascades, so an increase of the energy proportion of the hadronic cascades will enhance the muon production. Therefore, it is necessary to focus on reducing the production of pion since $1/3$ of its energy will transfer to electromagnetic cascade through $\pi^{0}\rightarrow 2\gamma$ decay. Kaons (such as $K^{+}$, $K^{-}$, $K^{0}_{S}$, $K^{0}_{L}$) participate in hadronic interactions, as in Eq.8, after reaching the critical energy, their decay products include not only muons but also pions. The branching ratio of kaon decaying into $\pi^{0}$ is much lower than decaying into muon, as reported in Patrignani et al. (2016). Thus, in this view, increasing the production of secondary particle kaons can increase the number of muons in the simulation after p-p collisions. Since kaons consist of a strange quark and a light quark, increasing the production of strange quarks is a key factor considered in this study.